OpenStudy (anonymous):

2x – 2y = –12 4x – y = 3

OpenStudy (anonymous):

what are you trying to find?

OpenStudy (anonymous):

When using the substitution method, what expression for y in the second equation would be substituted in the first equation?

OpenStudy (anonymous):

Well usually for substitution, you're given a value to start with. Since you don't, solve for y by solving the system through elimination.

OpenStudy (anonymous):

2x - 2y = -12 4x - y = 3 ------------------ 2x - 2y = -12 y = 4x - 3 ------------------ 2x - 2(4x - 3) = -12 2x - 8x - 6 = -12 -6x = -6 x = 1

OpenStudy (anonymous):

y = 5

OpenStudy (texaschic101):

4x - y = 3 -y = 3 - 4x (multiply by -1 to make y positive) y = 4x - 3 (this can be substituted in for y in the first equation)